# DisCoRD: Discrete Tokens to Continuous Motion via Rectified Flow Decoding

### v3

@@ Abstract
Human motion is inherently continuous and dynamic, posing significant challenges for generative models. While discrete generation methods are widely used, they suffer from limited expressiveness and frame-wise noise artifacts. In contrast, continuous approaches produce smoother, more natural motion but often struggle to adhere to conditioning signals due to high-dimensional complexity and limited training data. To resolve this “discord” between discrete and continuous representations we introduce DisCoRD: Discrete Tokens to Continuous Motion via Rectified Flow Decoding, a novel method that leverages rectified flow to decode discrete motion tokens in the continuous, raw motion space. Our core idea is to frame token decoding as a conditional generation task, ensuring that DisCoRD captures fine-grained dynamics and achieves smoother, more natural motions. Compatible with any discrete-based framework, our method enhances naturalness without compromising faithfulness to the conditioning signals on diverse settings. Extensive evaluations demonstrate that DisCoRD achieves state-of-the-art performance, with FID of 0.032 on HumanML3D and 0.169 on KIT-ML. These results establish DisCoRD as a robust solution for bridging the divide between discrete efficiency and continuous realism. Code and checkpoints will be released.
@@ 3.1 Preliminaries
[…]
@@ 3.2 DisCoRD
[…]
@@ 4.2 Quantitative results
By selecting the rectified flow algorithm from among the various diffusion model variants, we exploit its efficient transport mechanism to achieve inference speeds comparable to baseline models. As shown in Figure 5, we evaluated the decoding times for both Momask and our model on tokens generated by a pretrained token generator. At the default setting of 16 sampling steps, our model achieves decoding speeds on par with MoMask while delivering superior sJPE, stage 1 FID, and stage 2 FID. Furthermore, by reducing the sampling steps, our method can decode tokens significantly fasterer than MoMask, maintaining comparable or enhanced FID and sJPE performance. Additionally, although it was not the primary focus of this experiment, we observed that sJPE responds more sensitively to changes in sampling steps compared to FID, further confirming that our sJPE metric effectively captures subtle variations in motion quality.
@@
Table 5: Ablation studies. Evaluation on the HumanML3D test set assessing the impact of decoding strategies, projection methods, and training strategies. FF and RF denote feedforward and rectified flow model, respectively.
@@ 4.3 Qualitative Results
User Studies.
We conduct two user studies to (1) validate our motivation and method effectiveness and (2) evaluate how well sJPE aligns with human perception. The first study, shown in Figure 7, indicates that the discrete model Momask outperforms the continuous model MDM in faithfulness but lags in naturalness. In contrast, DisCoRD surpasses both, demonstrating its ability to generate motion that is both natural and faithful. In the second study, we find that sJPE exhibits 2.7 times higher correlation with human preference for naturalness compared to MPJPE, highlighting its effectiveness in evaluating sample-wise motion naturalness. Details of user studies are in Supplementary Section C.4 and E.2.
Visualizations.
We visualize motion trajectories in Figure 4, where DisCoRD, unlike discrete methods, produces smooth and expressive motion with low sJPE. Additionally, as shown in Figure 6, our method generates motions samples that closely align with textual descriptions while preserving a high degree of naturalness. Extensive visualizations are in Supplementary Section C.3 and E.1.
Figure 6: Qualitative comparisons on the test set of HumanML3D.
Figure 7: User study results on the HumanML3D dataset. Each bar represents a comparison between two models, with win rates depicted in blue and loss rates in red, evaluated based on naturalness and faithfulness.

### v4

@@ Abstract
Human motion is inherently continuous and dynamic, posing significant challenges for generative models. While discrete generation methods are widely used, they suffer from limited expressiveness and frame-wise noise artifacts. In contrast, continuous approaches produce smoother, more natural motion but often struggle to adhere to conditioning signals due to high-dimensional complexity and limited training data. To resolve this “discord” between discrete and continuous representations we introduce DisCoRD: Discrete Tokens to Continuous Motion via Rectified Flow Decoding, a novel method that leverages rectified flow to decode discrete motion tokens in the continuous, raw motion space. Our core idea is to frame token decoding as a conditional generation task, ensuring that DisCoRD captures fine-grained dynamics and achieves smoother, more natural motions. Compatible with any discrete-based framework, our method enhances naturalness without compromising faithfulness to the conditioning signals on diverse settings. Extensive evaluations demonstrate that DisCoRD achieves state-of-the-art performance, with FID of 0.032 on HumanML3D and 0.169 on KIT-ML. These results establish DisCoRD as a robust solution for bridging the divide between discrete efficiency and continuous realism. Project website: https://whwjdqls.github.io/discord-motion/
@@ 3.1 Preliminaries
$\begin{gathered}\min_{v}\int_{0}^{1}\mathbb{E}\left[\left\|(x_{1}-x_{0})-v(x_{t},t)\right\|^{2}\right]\,dt.\end{gathered}$ (3)
@@ 3.2 DisCoRD
$\begin{gathered}\min_{v}\int_{0}^{1}\mathbb{E}\left[\left\|(\mathbf{X}_{1}-\mathbf{X}_{0})-v(\mathbf{X}_{t},t,\mathbf{C})\right\|^{2}\right]\,dt,\\ \text{with}\quad\mathbf{X}_{t}=t\mathbf{X}_{1}+(1-t)\mathbf{X}_{0}.\end{gathered}$ (4)
@@ 4.2 Quantitative results
By selecting the rectified flow algorithm from among the various diffusion model variants, we exploit its efficient transport mechanism to achieve inference speeds comparable to baseline models. As shown in Figure 5, we evaluated the decoding times for both Momask and our model on tokens generated by a pretrained token generator. At the default setting of 16 sampling steps, our model achieves decoding speeds on par with MoMask while delivering superior sJPE, stage 1 FID, and stage 2 FID. Furthermore, by reducing the sampling steps, our method can decode tokens significantly fasterer than MoMask, maintaining comparable or enhanced FID and sJPE performance. Additionally, although it was not the primary focus of this experiment, we observed that sJPE responds more sensitively to changes in sampling steps compared to FID, further confirming that our sJPE metric effectively captures subtle variations in motion quality. Detailed decoding times are reported in the Supplementary Table E.
Effect of Residual Quantization.
We evaluate the impact of the representational capacity of motion tokens on DisCoRD by varying the residual quantization (RQ) levels in MoMask. As shown in Tab. 5, while DisCoRD consistently improves FID across all RQ levels, its R-Precision gains over each MoMask baseline increase with higher RQ levels, suggesting that DisCoRD better leverages richer token representations. This result is consistent with Table 2 where R-Precision drops for T2M-GPT (based on standard VQ-VAE) but maintains performance for MoMask and BAMM.
Methods Recon Generation
[…]
MoMask‑R0 0.125 0.082 0.505 0.696 0.793
+Ours 0.061(+51%) 0.068(+17%) 0.491(-2.8%) 0.682(-2.0%) 0.782(-1.4%)  00
MoMask‑R1 0.066 0.058 0.517 0.710 0.805
+Ours 0.024(+64%) 0.037(+36%) 0.510(-1.4%) 0.703(-1.0%) 0.800(-0.6%)  00
MoMask‑R2 0.046 0.053 0.520 0.712 0.806
+Ours 0.018(+61%) 0.031(+42%) 0.517(-0.6%) 0.709(-0.4%) 0.806(0.0%)
MoMask‑R4 0.024 0.047 0.519 0.711 0.807
+Ours 0.011(+54%) 0.032(+32%) 0.520(+0.2%) 0.711(0.0%) 0.806(-0.1%)  00
MoMask‑R5 0.019 0.045 0.521 0.713 0.807
+Ours 0.011(+42%) 0.032(+29%) 0.524(+0.6%) 0.715(+0.3%) 0.809(+0.2%)  00
Table 5: Effect of residual quantization levels (R). in MoMask on DisCoRD performance. Higher R yields greater R-Precision (RP) gains, indicating better use of richer representations.
@@
Table 6: Ablation studies. Evaluation on the HumanML3D test set assessing the impact of decoding strategies, projection methods, and training strategies. $F$ and RF denote feedforward and rectified flow model, respectively.
@@ 4.3 Qualitative Results
We visualize motion trajectories in Figure 4, where DisCoRD, unlike discrete methods, produces smooth and expressive motion with low sJPE. Extensive visualizations are provided in Supplementary Sections C.3 and E.1 with supporting user studies in Sections C.4 and E.2.
@@ 5 Conclusion and Discussion
Acknowledgments.
This study was supported by the Institute of Information & Communications Technology Planning & Evaluation (IITP) grant funded by the Korea government (MSIT) (No. RS-2024-00457882, Artificial Intelligence Research Hub Project). This work was also supported by the Culture, Sports and Tourism R&D Program through the Korea Creative Content Agency, funded by the Ministry of Culture, Sports and Tourism in 2024 (Project Name: Development of multimodal UX evaluation platform technology for XR spatial responsive content optimization, Project Number: RS-2024-00361757). T.-H. Oh was partially supported by the IITP grant funded by the Korea government (MSIT) (No. RS-2023-00225630, Development of Artificial Intelligence for Text-based 3D Movie Generation).
Supplementary Material
This supplementary material is organized as follows: Section A details the implementation of DisCoRD. Section B provides additional information on the datasets and evaluation metrics. Section C offers a comprehensive analysis of sJPE. Section D presents quantitative results excluded from the main paper. Section E includes additional qualitative results. We highly recommend viewing the accompanying video, as static images are insufficient to fully convey the intricacies of motion.
Appendix A Implementation Details
Table A provides an overview of the implementation details for our method. These configurations were employed to train the DisCoRD decoder using the pretrained Momask [12] quantizer. Specifically, the 512-dimensional codebook embeddings from MoMask are projected into the conditional channel dimension. This projection is concatenated with Gaussian noise of the same dimensionality as the output channel. The concatenated representation is subsequently projected into the input channel dimension of the U-Net architecture. The U-Net processes this input and transforms it back into the output channel dimension, generating the final output shape. For training, we used an input window size of 64 and trained the model for 35 hours on a single NVIDIA RTX 4090 Ti GPU.
Training Details
Optimizer AdamW $(0.9,0.999)$
LR 0.0005
LR Decay Ratio 0
LR Scheduler Cosine
Warmup Epochs 20
Gradient Clipping 1.0
Weight EMA 0.999
Flow Loss MSE Loss
Batch Size 768
Window Size 64
Steps 481896
Epochs 200
Model Details
Input Channels 512
Output Channels 263
Condition Channels 256
Activation SiLU
Dropout 0
Width (512, 1024)
# Resnet / Block 2
# Params 66.9M
Table A: Implementation details for training the DisCoRD decoder on the HumanML3D dataset using the pretrained Momask quantizer.
Appendix B Datasets and Evaluations
In this section, we provide additional explanations regarding the co-speech gesture generation and music-to-dance generation tasks that we were unable to describe in detail in the main paper.
B.1 Datasets.
For the co-speech gesture generation task, we utilized the SHOW dataset [62], a 3D holistic body dataset comprising 26.9 hours of in-the-wild talking videos. For the music-driven dance generation task, we used a mixed dataset combining AIST++ [25] and HumanML3D [11] where AIST++ is a large-scale 3D dance dataset created from multi-camera videos accompanied by music of varying styles and tempos, containing 992 high-quality pose sequences in the SMPL format.
B.2 Evaluations.
To evaluate the co-speech gesture generation task, we used Frechet Gesture Distance (FGD) [33], which measures the difference between the latent distributions of generated and real motions. Since our focus is on body movements, we reported body FGD, which quantifies differences specifically for the body part, for ProbTalk [30]. For TalkSHOW [62], which only utilizes holistic FGD—a metric that measures differences across the entire motion, including the face and hands—we reported the holistic FGD. To evaluate the music-to-dance generation task, we utilized $\mathrm{Dist_{k}}$, which quantifies the distributional spread of generated dances based on kinetic features, and $\mathrm{Dist_{g}}$, which does the same for geometric features, as proposed in [53]. A smaller difference between the distributions of the generated motion and the ground truth motion indicates that the $\mathrm{Dist_{k}}$ and $\mathrm{Dist_{g}}$ values of the generated motion align closely with those of the ground truth, reflecting a similar level of distributional spread.
Appendix C Additional Analysis on sJPE
To evaluate the sample-wise naturalness of reconstructed motions, we introduce the symmetric Jerk Percentage Error (sJPE), as defined in Equation 5 of the main paper. We present detailed formulations of Noise sJPE and Static sJPE, supported by analysis using generated motion samples. Furthermore, qualitative comparisons highlight the effectiveness of DisCoRD against state-of-the-art discrete methods. Finally, we investigate the alignment of sJPE with human preference to validate its perceptual relevance.
C.1 Visualization of Fine-Grained Motion
To analyze fine-grained motion trajectories, we follow a three-step procedure. First, we select a joint for visualization, typically hand joints due to their high dynamism, and track their positional changes over time. Second, we apply a Gaussian filter to smooth the trajectory, reducing noise. Finally, we compute the difference between the smoothed and original trajectories to isolate fine-grained motion components. This method allows for detailed evaluation of frame-wise noise and under-reconstructed regions in motion trajectories. The visualizations in Figure 5 of the main paper and the qualitative samples in the supplementary material are generated using this process.
C.2 Details on sJPE.
Within the symmetric Jerk Percentage Error (sJPE), we define two components: Noise sJPE and Static sJPE. These isolate the instances where the predicted jerk overestimates or underestimates the ground truth jerk, respectively.
Noise sJPE and Static sJPE.
Noise sJPE captures the average overestimation of jerk in the predicted motion signal, meaning frame-wise noise, corresponding to cases where $J_{\text{pred},t}>J_{\text{true},t}$. It is defined as:
$\text{Noise sJPE}=\frac{1}{n}\sum_{t=1}^{n}\frac{\max\left(0,J_{\text{pred},t}-J_{\text{true},t}\right)}{|J_{\text{true},t}|+|J_{\text{pred},t}|}.$ (6)
The operator $\max(0,x)$ ensures that only positive differences contribute to Noise sJPE, separating overestimations from underestimations.
Noise sJPE can be seen on the red box of Figure A. Time steps where motion trajectory is noisy compared to ground truth motion show bigger jerk. The area under the predicted jerk and above the ground truth jerk, shown in blue area, is proportional to the Noise sJPE, meaning frame-wise noise.
Static sJPE measures the average underestimation of jerk, meaning lack of dynamism in the predicted motion, corresponding to cases where $J_{\text{pred},t}\leq J_{\text{true},t}$. It is defined as:
$\text{Static sJPE}=\frac{1}{n}\sum_{t=1}^{n}\frac{\max\left(0,J_{\text{true},t}-J_{\text{pred},t}\right)}{|J_{\text{true},t}|+|J_{\text{pred},t}|}.$ (7)
Static sJPE can be seen on the green box of Figure A. Time steps where motion trajectory is under-reconstructed compared to ground truth motion show smaller jerk. The area above the predicted jerk and under the ground truth jerk, shown in red area, is proportional to the Static sJPE, meaning under reconstructed motions.
The overall sJPE can be expressed as the sum of Noise sJPE and Static sJPE:
[…]
These formulations provide a measure of prediction accuracy by separately accounting for the tendencies of the predictive model to overestimate or underestimate the true motion jerks.
Figure A: Relationship between fine-grained trajectory and jerk: Frame-wise noise in predicted motions, highlighted in the red box, results in higher jerk values compared to the ground truth, represented by the blue areas. The sum of the blue areas corresponds to Noise sJPE. Conversely, under-reconstruction in predicted motions, highlighted in the green box, leads to lower jerk values compared to the ground truth, represented by the red areas. The sum of the red areas corresponds to Static sJPE.
Figure B: Joint Trajectory and Jerk: Under-Reconstruction in Discrete Methods DisCoRD effectively reduces the red area, demonstrating its capability to reconstruct dynamic motion accurately. This improvement is also reflected in the lower sJPE value.
Figure C: Joint Trajectory and Jerk: Frame-Wise Noise in Discrete Methods DisCoRD significantly reduces the blue area, indicating its ability to generate smooth motions that closely resemble the ground truth. This improvement is further reflected in the lower sJPE value.
C.3 Qualitative Results on Joint Trajectory and Jerk
We present a series of figures demonstrating the effectiveness of DisCoRD in reconstructing smooth and dynamic motion. For each sample, the first row visualizes the motion trajectory, while the second row plots the corresponding jerk at each time step, with the calculated sJPE displayed at the top. This visualization enables detailed analysis of fine-grained trajectories in predicted motions and highlights the contributions of Noise sJPE and Static sJPE to the overall sJPE.
We compare DisCoRD with recent discrete methods, including T2M-GPT [64], MMM [41], and Momask [12]. Motion samples that illustrate under-reconstruction in discrete methods are presented in Figure B, while those that exhibit frame-wise noise are shown in Figure C. Samples showing both issues in discrete models are displayed in Figure D. DisCoRD effectively reduces frame-wise noise while accurately reconstructing dynamic, fast-paced motions. This is shown in both the visualizations and the sJPE results.
Figure D: Joint Trajectory and Jerk: Both Frame-Wise Noise and Under-Reconstruction in Discrete Methods DisCoRD addresses both frame-wise noise and under-reconstruction by simultaneously reducing the blue and red areas. This demonstrates its ability to generate smooth and dynamic motions, closely aligning with the ground truth. This further supported by the lower sJPE values.
C.4 Correlation between sJPE and human perception.
To further verify that sJPE aligns with human judgment of naturalness, we conducted an additional user study. We asked participants to rank three models—MLD, MoMask, and ours—in order of naturalness, guided as Figure J. The user interface for this user study is shown in Figure L. The rankings were scored such that the first place received 1 point, the second place 2 points, and the third place 3 points. Using these human scores, we calculated Pearson’s correlation between the human scores and two metrics—MPJPE and sJPE— for each sample. During this process, we excluded the lowest 10% of samples in terms of human score standard deviation among models, as these were considered indistinguishable by human evaluators. Our analysis revealed that the average Pearson’s correlation between MPJPE and human scores was 0.181, whereas the correlation between sJPE and human scores was significantly higher at 0.483. This result demonstrates the effectiveness of sJPE in evaluating sample-wise naturalness.
Appendix D Additional Quantitative Results
D.1 Performance on Text-to-Motion Generation.
In Table B, we present a comparison of our method against additional results from various text-to-motion models. Our method consistently achieves strong performance on the HumanML3D and KIT-ML [42] test sets, even when evaluated alongside these additional models. While ReMoDiffuse achieves particularly strong performance on KIT-ML, it is worth noting that its performance benefits from the use of a specialized database for high-quality motion generation, which makes direct comparisons less appropriate.
Datasets Methods R Precision $\uparrow$ FID $\downarrow$ MultiModal Dist $\downarrow$ MultiModality $\uparrow$
Top 1 Top 2 Top 3
Human ML3D MDM [52] - - ${0.611}^{\pm{.007}}$ ${0.544}^{\pm{.044}}$ ${5.566}^{\pm{.027}}$ $\underline{{2.799}}^{\pm{.072}}$
MLD [6] ${0.481}^{\pm{.003}}$ ${0.673}^{\pm{.003}}$ ${0.772}^{\pm{.002}}$ ${0.473}^{\pm{.013}}$ ${3.196}^{\pm{.010}}$ ${2.413}^{\pm{.079}}$
Fg-T2M [wang2023fg] ${0.492}^{\pm{.002}}$ ${0.683}^{\pm{.003}}$ ${0.783}^{\pm{.024}}$ ${0.243}^{\pm{.019}}$ ${3.109}^{\pm{.007}}$ ${1.614}^{\pm{.049}}$
M2DM [m2dm] ${0.497}^{\pm{.003}}$ ${0.682}^{\pm{.002}}$ ${0.763}^{\pm{.003}}$ ${0.352}^{\pm{.005}}$ ${3.134}^{\pm{.010}}$ $\mathbf{{3.587}}^{\pm{.072}}$
M2D2M [7] - - ${0.799}^{\pm{.002}}$ ${0.087}^{\pm{.004}}$ ${3.018}^{\pm{.010}}$ ${2.115}^{\pm{.079}}$
MotionGPT [zhang2024motiongpt] ${0.364}^{\pm{.005}}$ ${0.533}^{\pm{.003}}$ ${0.629}^{\pm{.004}}$ ${0.805}^{\pm{.002}}$ ${3.914}^{\pm{.013}}$ ${2.473}^{\pm{.041}}$
MotionLLM [59] ${0.482}^{\pm{.004}}$ ${0.672}^{\pm{.003}}$ ${0.770}^{\pm{.002}}$ ${0.491}^{\pm{.019}}$ ${3.138}^{\pm{.010}}$ -
MotionGPT-2 [motiongpt2] ${0.496}^{\pm{.002}}$ ${0.691}^{\pm{.003}}$ ${0.782}^{\pm{.004}}$ ${0.191}^{\pm{.004}}$ ${3.080}^{\pm{.013}}$ ${2.137}^{\pm{.022}}$
AttT2M [zhong2023attt2m] ${0.499}^{\pm{.003}}$ ${0.690}^{\pm{.002}}$ ${0.786}^{\pm{.002}}$ ${0.112}^{\pm{.006}}$ ${3.038}^{\pm{.007}}$ ${2.452}^{\pm{.051}}$
+ DisCoRD (Ours) ${0.476}^{\pm{.008}}$ ${0.663}^{\pm{.006}}$ ${0.760}^{\pm{.007}}$ ${0.095}^{\pm{.011}}$ ${3.121}^{\pm{.009}}$ ${1.831}^{\pm{.048}}$
+ DisCoRD (Ours) ${0.522}^{\pm{.003}}$ $\underline{{0.715}}^{\pm{.005}}$ $\underline{{0.811}}^{\pm{.004}}$ $\underline{{0.041}}^{\pm{.002}}$ $\underline{{2.921}}^{\pm{.015}}$ ${1.772}^{\pm{.067}}$
+ DisCoRD (Ours) $\underline{{0.524}}^{\pm{.003}}$ $\underline{{0.715}}^{\pm{.003}}$ ${0.809}^{\pm{.002}}$ $\mathbf{{0.032}}^{\pm{.002}}$ ${2.938}^{\pm{.010}}$ ${1.288}^{\pm{.043}}$
MLD [6] ${0.390}^{\pm{.008}}$ ${0.609}^{\pm{.008}}$ ${0.734}^{\pm{.007}}$ ${0.404}^{\pm{.027}}$ ${3.204}^{\pm{.027}}$ ${2.192}^{\pm{.071}}$
Fg-T2M [wang2023fg] ${0.418}^{\pm{.005}}$ ${0.626}^{\pm{.004}}$ ${0.745}^{\pm{.004}}$ ${0.571}^{\pm{.047}}$ ${3.114}^{\pm{.015}}$ ${1.019}^{\pm{.029}}$
M2DM [m2dm] ${0.416}^{\pm{.004}}$ ${0.628}^{\pm{.004}}$ ${0.743}^{\pm{.004}}$ ${0.515}^{\pm{.029}}$ ${3.015}^{\pm{.017}}$ $\mathbf{{3.325}}^{\pm{.037}}$
M2D2M [7] - - ${0.753}^{\pm{.006}}$ ${0.378}^{\pm{.023}}$ ${3.012}^{\pm{.021}}$ ${2.061}^{\pm{.067}}$
MotionGPT [zhang2024motiongpt] ${0.340}^{\pm{.002}}$ ${0.570}^{\pm{.003}}$ ${0.660}^{\pm{.004}}$ ${0.868}^{\pm{.032}}$ ${3.721}^{\pm{.018}}$ ${2.296}^{\pm{.022}}$
MotionLLM [59] ${0.409}^{\pm{.006}}$ ${0.624}^{\pm{.007}}$ ${0.750}^{\pm{.005}}$ ${0.781}^{\pm{.026}}$ ${2.982}^{\pm{.022}}$ -
MotionGPT-2 [motiongpt2] ${0.427}^{\pm{.003}}$ ${0.627}^{\pm{.002}}$ ${0.764}^{\pm{.003}}$ ${0.614}^{\pm{.005}}$ ${3.164}^{\pm{.013}}$ $\underline{{2.357}}^{\pm{.022}}$
AttT2M [zhong2023attt2m] ${0.413}^{\pm{.006}}$ ${0.632}^{\pm{.006}}$ ${0.751}^{\pm{.006}}$ ${0.870}^{\pm{.039}}$ ${3.039}^{\pm{.021}}$ ${2.281}^{\pm{.047}}$
+ DisCoRD (Ours) ${0.382}^{\pm{.007}}$ ${0.590}^{\pm{.007}}$ ${0.715}^{\pm{.004}}$ ${0.541}^{\pm{.038}}$ ${3.260}^{\pm{.028}}$ ${1.928}^{\pm{.059}}$
+ DisCoRD (Ours) $\mathbf{{0.434}}^{\pm{.007}}$ $\mathbf{{0.657}}^{\pm{.005}}$ $\underline{{0.775}}^{\pm{.004}}$ $\underline{{0.169}}^{\pm{.010}}$ $\underline{{2.792}}^{\pm{.015}}$ ${1.266}^{\pm{.046}}$
Table B: Additional quantitative evaluation on the HumanML3D and KIT-ML test sets. $\pm$ indicates a 95% confidence interval. +DisCoRD indicates that the baseline model’s decoder is replaced with our DisCoRD decoder. Bold indicates the best result, while underscore refers the second best.
D.2 Performance on Various Tasks.
In Table C, we present additional evaluation results for co-speech gesture generation. Following [62], we additionally report Diversity, which measures the variance among multiple samples generated from the same condition, and Beat Consistency (BC), which evaluates the synchronization between the generated motion and the corresponding audio. In Table D, we provide additional evaluation results for music-to-dance generation. Following [46], we report $\mathrm{FID_{k}}$ to measure differences in kinetic motion features and $\mathrm{FID_{g}}$ for geometric motion features. Additionally, we include the Beat Align Score (BAS) to assess the synchronization between motion and music. While  [53] has shown that these metrics are not fully reliable and often fail to align with actual output quality, we include them to follow established conventions.
Methods Diversity $\uparrow$ BC $\rightarrow$ (0.868)
TalkSHOW [62] 0.821 0.872
+DisCoRD(Ours) 0.919 0.876
ProbTalk [30] 0.259 0.795
+DisCoRD(Ours) 0.331 0.866
Table C: Additional quantitative results on each method’s SHOW test set. The results demonstrate that our method performs on par with, or surpasses, the baseline models.
Methods $\mathrm{FID_{k}}\downarrow$ $\mathrm{FID_{g}}\downarrow$ BAS $\uparrow$
Ground Truth 17.10 10.60 0.2374
TM2D [10] 19.01 20.09 0.2049
+DisCoRD(Ours) 23.98 88.74 0.2190
Table D: Additional quantitative results on the AIST++ test set. The results demonstrate that, although our method shows performance degradation on $\mathrm{FID_{k}}$ and $\mathrm{FID_{g}}$, which are known to be unreliable, it achieves improvement in the Beat Align Score.
Appendix E Additional Qualitative Results
E.1 Motion visualization.
In Figure F and Figure G, we present qualitative comparisons between our model and other leading approaches. In Figure H, we additionally display more qualitative results of our method. We observed that our method effectively follows the text prompts while maintaining naturalness in the generated outputs. Again, we highly recommend viewing the accompanying video, as static images are insufficient to fully convey the intricacies of motion.
Figure E: User study results on the HumanML3D dataset. Each bar represents a comparison between two models, with win rates depicted in blue and loss rates in red, evaluated based on naturalness and faithfulness.
E.2 User preference study details.
We conduct two user studies to (1) validate our motivation and method effectiveness and (2) evaluate how well sJPE aligns with human perception. The first study, shown in Figure E, indicates that the discrete model Momask outperforms the continuous model MDM in faithfulness but lags in naturalness. In contrast, DisCoRD surpasses both, demonstrating its ability to generate motion that is both natural and faithful. In the second study, we find that sJPE exhibits 2.7 times higher correlation with human preference for naturalness compared to MPJPE, highlighting its effectiveness in evaluating sample-wise motion naturalness. Participants were guided to evaluate both faithfulness and naturalness, as shown in Figure I. Given two motion videos generated by two different models on the same prompt, participants were asked to choose a better one in terms of faithfulness and naturalness, as shown in Figure K. Total 41 participants participated in this user study.
Figure F: Qualitative comparisons on the test set of HumanML3D.
Figure G: Additional qualitative comparisons on the HumanML3D test set. The continuous method, MLD, often fails to perfectly align with the text consistently, while the discrete method, MoMask, exhibits issues such as under-reconstruction, resulting in minimal hand movement, or unnatural leg jitter caused by frame-wise noise.
Figure H: Additional qualitative results of our method on the HumanML3D test set.
Method Sampling Steps
2 16 50 100
DDPM (DDIM sample) 0.055 / 0.007s 0.044 / 0.087s 0.038 / 0.331s 0.035 / 0.689s
Linear SDE 8.998 / 0.024s 0.047 / 0.157s 0.033 / 0.472s 0.032 / 0.955s
Ours 0.034 / 0.016s 0.032 / 0.221s 0.032 / 0.703s 0.032 / 1.426s
Table E: FID and decoding time (s) for different sampling steps $(\downarrow)$. The original MoMask has a decoding time of 0.244 seconds. We trained a diffusion decoder (DDPM) using the same architecture as our rectified flow decoder for a fair comparison. For the Linear SDE variant, we replaced only the sampler in our model with the Euler-Maruyama sampler, keeping all other components identical.
Figure I: Guidelines for user study in the Main paper: participants were asked to evaluate Faithfulness and Naturalness, excluding hand and facial movements that are not included in HumanML3D.
Figure J: Guidelines for User Study in the Supplementary: Participants were asked to evaluate Naturalness, excluding hand and facial movements that are not included in HumanML3D.
Figure K: User evaluation interface for the user study in the Main paper: participants were presented with two randomly selected videos and asked to choose the better sample in terms of faithfulness and naturalness.
Figure L: User evaluation interface for the user study in the supplementary: participants were presented with a grid layout containing the GT video and three generated videos. Using the GT video as the upper bound, they were asked to rank the three generated videos in terms of naturalness.